OpenStudy (nick1232):

-28= −2( −r+8)+2(3+8r)

OpenStudy (nick1232):

please help

OpenStudy (anonymous):

-28 = 2r - 16 + 6 + 16r -28 = -10 + 18r -18 = 18r R = -1

OpenStudy (nick1232):

ok thank you